# Joint Pricing and Resource Allocation: An Optimal Online-Learning Approach

### v1

@@ 7 Conclusion
Appendix A Proof Details
A.1 Proof of Theorem 3.1
Proof.
We conduct the proof according to the proof sketch displayed after Theorem 3.1 in the main paper.
Firstly, we construct the dual problem of Eq. 2. Here we denote $\tilde{C}_{i,j}(p):=-p+C_{i,j}$, and we know that the sorting of $\tilde{C}_{i,j}$ is identical to that of $C_{i,j}$ for any $p\in\mathbb{R}$. The Lagrangian can be written as
$L(X,A,\mu,\lambda):=$ $\sum_{i=1}^{m}\sum_{j=1}^{n}\tilde{C}_{i,j}X_{i,j}-\sum_{i=1}^{m}\sum_{j=1}^{n}A_{i,j}X_{i,j}+\sum_{j=1}^{n}\mu_{j}(\sum_{i=1}^{m}X_{i,j}-D_{j})+\sum_{i=1}^{m}\lambda_{i}(\sum_{j=1}^{n}X_{i,j}-I_{i})$ (11)
$=$ $\sum_{i}\sum_{j}(\tilde{C}_{i,j}-A_{i,j}+\mu_{j}+\lambda_{i})X_{i,j}-\sum_{j}\mu_{j}D_{j}-\sum_{i}\lambda_{i}I_{i}.$
Here $A\in\mathbb{R}^{m\times n}_{+},\mu\in\mathbb{R}^{n}_{+},\lambda\in\mathbb{R}^{m}_{+}$. Therefore, the primal problem can be written as $\min_{X}\max_{A,\mu,\lambda}L(X,A,\mu,\lambda)$, and the dual problem is:
$\max_{A,\mu,\lambda}\min_{X}L(X,A,\mu,\lambda)$ (12)
[…]
This comes from the complementary slackness. Since $\tilde{C}_{i_{k},j_{k}}<0$ for $k\leq K$, the condition of $\tilde{C}_{i_{k},j_{k}}-A_{i_{k},j_{k}}+\mu_{j_{k}}+\lambda_{i_{k}}=0$ is not necessarily satisfiable for those $k\in[K]$. Given this, the dual problem can be equivalently written as
$-\min_{\mu,\lambda}$ $\sum_{j=1}^{n}\mu_{j}D_{j}+\sum_{i=1}^{m}\lambda_{i}I_{i}$ (13)
s.t. $\tilde{C}_{i_{k},j_{k}}+\mu_{j_{k}}+\lambda_{i_{k}}\geq 0,\forall k\leq K.$
Denote $h(\mu,\lambda;\vec{I},\vec{D}):=\sum_{j=1}^{n}\mu_{j}D_{j}+\sum_{i=1}^{m}\lambda_{i}I_{i}$. Also, define a new equation
$\tilde{h}(\mu,\lambda;\tilde{\vec{I}},\tilde{\vec{D}},\vec{Q}):=\sum_{j=1}^{n}\mu_{j}\tilde{D}_{j}+\sum_{i=1}^{m}\lambda_{i}\tilde{I}_{i}+\sum_{k=1}^{mn}Q_{k}(\lambda_{i_{k}}+\mu_{j_{k}}).$ (14)
By initializing $\vec{Q}=0$ and $\tilde{\vec{I}}=\vec{I},\tilde{\vec{D}}=\vec{D}$, we know that $h(\mu,\lambda;\vec{I},\vec{D})=\tilde{h}(\mu,\lambda;\tilde{\vec{I}},\tilde{\vec{D}},\vec{Q})$. In the following, we conduct a series of equivalent transformations on $\tilde{h}$, which will further induce an algorithm to solve the dual problem.
Algorithm 8 Equivalent transformation on $\tilde{h}$
  Initialization: $\vec{Q}=0,\tilde{\vec{I}}=\vec{I},\tilde{\vec{D}}=\vec{D}$.
  for $k=1,2,\ldots,K$, do
     Set $Q_{k}\leftarrow\min\{\tilde{D}_{j_{k}},\tilde{I}_{i_{k}}\}$.
     Update $\tilde{I}_{i_{k}}\leftarrow\tilde{I}_{i_{k}}-Q_{k}$.
     Update $\tilde{D}_{j_{k}}\leftarrow\tilde{D}_{j_{k}}-Q_{k}$.
  end for
Notice that
  1. i
    The value of $\tilde{h}(\mu,\lambda;\tilde{\vec{I}},\tilde{\vec{D}},\vec{Q})$ is unchanged over the processing of Algorithm 8.
  2. ii
    The $\{Q_{k}\}_{k=1}^{mn}$ series we get from Algorithm 8 is identical to the $\{X_{i_{k},j_{k}}\}_{k=1}^{mn}$ series from Algorithm 1.
In the following, we focus on proving this lemma
Lemma A.1 (Feasible dual solutions).
The equation set of $\lambda\in\mathbb{R}_{+}^{m},\mu\in\mathbb{R}_{+}^{n}$
$&\mu_{j_{k}}+\lambda_{i_{k}}=-\tilde{C}_{i_{k},j_{k}},\qquad\qquad\text{ for }k\in\{k\leq K,Q_{k}\neq 0\}\\ &\mu_{j_{k}}=\lambda_{i_{k}}=0,\qquad\qquad\qquad\qquad\text{ otherwise.}\end{aligned}\right.$ (15)
has a feasible solution.
Before proving Lemma A.1, we firstly propose and prove the following series of lemma.
Lemma A.2.
After the operations of Algorithm 8, the total number of $k$ such that $Q_{k}\neq 0$ does not exceed $\min\{K,m+n\}$.
Proof of Lemma A.2.
According to Algorithm 8, the operation at time $k$ leads to either $\tilde{I}_{i_{k}}=0$ or $\tilde{D}_{j_{k}}=0$. On the one hand, there are at most $k$ times of operations. On the other hand, there are at most $(m+n)$ number of $\tilde{I}_{j}$ and $\tilde{D}_{j}$ that can be reduced to 00. ∎
Lemma A.3.
For $k_{1},k_{2}\in[K],k_{1}<k_{2}$, s.t. $Q_{k_{1}}\neq 0,Q_{k_{2}}\neq 0,i_{k_{1}}=i_{k_{2}}$, then for any $k>k_{1}$ we have either $Q_{k}=0$ or $j_{k}\neq j_{k_{1}}$.
Similarly, for $k_{3},k_{4}\in[k],k_{3}<k_{4}$, s.t. $Q_{k_{3}}\neq 0,Q_{k_{4}}\neq 0,j_{k_{3}}=j_{k_{4}}$, then for any $k>k_{3}$ we have either $Q_{k_{3}}=0$ or $i_{k}\neq i_{k_{3}}$.
Proof of Lemma A.3.
We only show the first case here, and the second case is by symmetry. Consider $\tilde{I}_{i_{k_{2}}}$ by the start of period $k=k_{2}$: Since $Q_{k_{2}}\neq 0$, we have $\tilde{I}_{i_{k_{2}}}>0$ at that time. Also, since $k_{1}<k_{2}$, we have $\tilde{I}_{i_{k_{1}}}>0$ by the end of period $k=k_{1}$. As a result, we have $\tilde{D}_{j_{k_{1}}}=0$ by the end of period $k=k_{1}$. Therefore, for any $k>k_{1}$ such that $j_{k}=j_{k_{1}}$, we have $\tilde{D}_{j_{k}}=0$ and therefore $Q_{k}=0$. ∎
Lemma A.4.
Define a bipartite graph $G(V,E)$, where $V=V_{1}\cup V_{2},V_{1}=[m],V_{2}=[n]$ and $E:=\{(i_{k},j_{k})\text{ for all }k\text{ s.t.}Q_{k}\neq 0\}$. There exists a series of independent trees $\{T_{l}\}_{l=1}^{t}$ such that $G=T_{1}\cup T_{2}\cup\ldots\cup T_{t}$. Here two independent trees $T_{i},T_{j}\subset G$ satisfies $T_{i}\cap T_{j}=\emptyset$.
Proof of Lemma A.4.
We sufficiently show that there exists no loop in this bipartite graph $G(V,E)$. In fact, according to Lemma A.3, for any edge $(i_{k_{1}},j_{k_{1}})\in E$, either $i_{k_{1}}$ or $j_{k_{1}}$ is a leaf node. On the other hand, there exists no leaf node in any loop. Therefore, no loop exists in this graph. ∎
With the help of Lemma A.2, Lemma A.3, Lemma A.4, we may process to prove Lemma A.1.
Proof of Lemma A.1.
Based on Lemma A.4, the set of tuples $\{(i_{k},j_{k}),k\in[K]\text{ s.t. }Q_{k}\neq 0\}$ can be separated as a series of disjoint "tree" sets. Therefore, we only need to prove Lemma A.1 holds on a connected "tree" set. Denote this tree as $T:=G_{T}(V_{T},E_{T})$, where $t=|T|$ is the size (i.e. number of nodes) of tree $T$, $E_{T}:=\{(i_{k_{\tau}},j_{k_{\tau}}),\tau=1,2,\ldots,t,1\leq k_{1}\leq k_{2}\leq\ldots\leq k_{t}\leq K\}$, and $V_{T}:=\{i_{k_{\tau}}\}_{\tau=1}^{t}\cup\{j_{k_{\tau}}\}_{\tau=1}^{t}$. In the following, we show that there exists $\{\mu_{j_{k_{\tau}}}\}_{\tau=1}^{t}\geq 0,\{\lambda_{i_{k_{\tau}}}\}_{\tau=1}^{t}\geq 0$ such that $\mu_{j_{k_{\tau}}}+\lambda_{i_{k_{\tau}}}=-\tilde{C}_{i_{k_{\tau}},j_{k_{\tau}}}$ for $\tau=1,2,\ldots,t$.
We prove this proposition by induction on $t$ (denoted as the size of tree $T$).
  1. (i)
    When $t=1$, the equation set reduces to $\mu_{j_{k_{1}}}+\lambda_{i_{k_{1}}}=-\tilde{C}_{i_{k_{1}},j_{k_{1}}}$. A feasible solution is $\mu_{j_{k_{1}}}=-\tilde{C}_{i_{k_{1}},j_{k_{1}}}$ and $\lambda_{i_{k_{1}}}=0$.
  2. (ii)
    Assume that the proposition holds for $t\leq N$ for some positive integer $N$. In the following, we show that it also holds for $t=N+1$.
    Consider $(i_{k_{1}},j_{k_{1}})$ which is the edge in the first place. According to Lemma A.3, we show that either $i_{k_{1}}$ or $j_{k_{1}}$ is a leaf of tree $T$: In fact, if $i_{k_{1}}$ is not a leaf, then $\exists\tau_{1}\geq 2$ s.t. $i_{k_{\tau_{1}}}=i_{k_{1}}$ and $j_{k_{\tau_{1}}}\neq j_{k_{1}}$, and therefore $k_{\tau_{1}}\geq k_{1}$. According to Lemma A.3, for any $k>k_{1},Q_{k}\neq 0$, we have $j_{k}\neq j_{k_{1}}$. As a result, we have $j_{k_{\tau}}\neq j_{j_{1}}$ for any $\tau\in[t],\tau\neq 1$. Similarly, if $j_{k_{1}}$ is not a leaf, then $i_{k_{1}}$ must be a leaf. Without loss of generality, assume $j_{k_{1}}$ is a leaf, and therefore $i_{k_{1}}$ is not a leaf (otherwise $T$ is not connected).
    Denote $T^{\prime}:=G(V_{T}\setminus\{j_{k_{1}}\},E_{T}\setminus\{(i_{k_{1}},j_{k_{1}})\})$ as the rest of the tree $T$ excluding $j_{k_{1}}$, and we have $|T^{\prime}|=N$ and $i_{k_{1}}\in T^{\prime}$. According to the induction hypothesis that holds for $t=N$. the equation set defined on tree $T^{\prime}$ has a feasible solution. Asa result, there exists $\tau^{*}\geq 2,\tau^{*}\in[t]$ such that $i_{k_{1}}=i_{k_{\tau^{*}}}$. Notice that $\lambda_{i_{k_{1}}}=\lambda_{i_{k_{\tau^{*}}}}\leq-\tilde{C}_{i_{k_{\tau^{*}}},j_{k_{\tau^{*}}}}\leq-\tilde{C}_{i_{k_{1}},j_{k_{1}}}$. Therefore, we know that $-\tilde{C}_{i_{k_{1}},j_{k_{1}}}-\lambda_{i_{k_{\tau^{*}}}}\geq 0$. By letting $\mu_{j_{k_{1}}}=-\tilde{C}_{i_{k_{1}},j_{k_{1}}}-\lambda_{i_{k_{\tau^{*}}}}$ and combine it with the solutions derived by induction hypothesis, we have construct a feasible solution to the equation set as $t=N+1$.
By the principle of induction, we have proven Lemma A.1. ∎
Now we are able to get back to the main stream of the proof. Denote the solutions to Eq. 15 as $(\mu^{*},\lambda^{*})$, and we have
$h(\mu^{*},\lambda^{*};\vec{I},\vec{D})=$ $\tilde{h}(\mu^{*},\lambda^{*};\tilde{\vec{I}},\tilde{\vec{D}},\tilde{\vec{Q}})$ (16)
$=$ $0+0+\sum_{k=1}^{K}Q_{k}\cdot(-\tilde{C}_{i_{k},j_{k}})$
$=$ $-\sum_{k=1}^{K}X_{i_{k},j_{k}}\tilde{C}_{i_{k},j_{k}}.$
Therefore, we get a feasible solution to the dual problem shown in Eq. 13 whose value equals $-h(\mu^{*},\lambda^{*};\vec{I},\vec{D})=\sum_{k=1}^{K}X_{i_{k},j_{k}}\tilde{C}_{i_{k},j_{k}}$. This further equals the value of the primal solution to Eq. 2 proposed by Algorithm 1. Due to the duality, we know that the primal solution is optimal. ∎
A.2 Proof of Lemma 3.2
@@ Appendix A Proof Details
$g(\vec{I},\vec{D}):=$ $\min_{x}\max_{\mu,\lambda,\eta\succeq 0}L(x,\mu,\lambda,\eta;\vec{I},\vec{D})=$ $\max_{\mu,\lambda,\eta\succeq 0}\min_{x}L(x,\mu,\lambda,\eta;\vec{I},\vec{D})$ (19)
@@
A.3 Proof of Lemma 3.4
At the beginning, we define a few key quantities and notations. We denote $\tilde{C}_{i,j}(p):=-p+C_{i,j}$ as we did in Section A.1. Also, we adopt the notations $\tilde{I}_{i_{k}}$ and $\tilde{D}_{j_{k}}$ in Algorithm 1 as the remaining inventories and demands at each step $k$ of allocating $X_{i_{k},j_{k}}$. Moreover, we denote $\{X_{i_{k},j_{k}}^{*}\}_{k=1}^{K}$ as the optimal solution to Eq. 2 generated by Algorithm 1. Given Algorithm 1 and Theorem 3.1, we know that $g(\vec{I},p,\vec{D})$ can be written as the following form
$g(\vec{I},p,\vec{D})=\sum_{k=1}^{K}\tilde{C}_{i_{k},j_{k}}(p)\cdot\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}=\sum_{k=1}^{K}\tilde{C}_{i_{k},j_{k}}(p)\cdot X_{i_{k},j_{k}}^{*}.$ (20)
We firstly show that $Q_{t}(\vec{I},p)$ is Lipschitz, which is sufficient to show that $g(\vec{I},p,\vec{D})$ is Lipschitz. In fact, for any $\Delta\vec{I}\geq 0$, we have $g(\vec{I}+\Delta\vec{I},p,\vec{D})\geq g(\vec{I},p,\vec{D})-\Delta\vec{I}\cdot C_{i_{1},j_{1}}$ and $g(\vec{I}+\Delta\vec{I},p,\vec{D})\leq g(\vec{I},p,\vec{D})$. Also, for any $\Delta\vec{D}\geq 0$, we have $g(\vec{I},p,\vec{D}+\Delta\vec{D})\geq g(\vec{I},p,\vec{D})-\Delta\vec{I}\cdot C_{i_{1},j_{1}}$ and $g(\vec{I},p,\vec{D}+\Delta\vec{D})\leq g(\vec{I},p,\vec{D})$. Moreover, for any $\Delta p\geq 0$, we have $g(\vec{I},p+\Delta p,\vec{D})\geq g(\vec{I},p,\vec{D})-\Delta p\cdot\|\vec{I}\|_{1}$ and $g(\vec{I},p+\Delta p,\vec{D})\leq g(\vec{I},p,\vec{D})+\Delta p\cdot\|\vec{I}\|_{1}$.
Since $Q_{t}(\vec{I},p)=\langle\vec{\gamma},\vec{I}\rangle+g(\vec{I},p,\vec{D}_{t}(p))$ where $\vec{D}_{t}(p)=\vec{a}-\vec{b}\cdot p+\vec{N}_{t}$, we have
$\frac{\partial Q_{t}(\vec{I},p)}{\partial p}=$ $\sum_{k=1}^{K}\frac{\partial\tilde{C}_{i_{k},j_{k}}(p)}{\partial p}\cdot\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}+\sum_{k=1}^{K}\tilde{C}_{i_{k},j_{k}}\cdot\frac{\partial\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}}{\partial p}$ (21)
$=$ $\sum_{k=1}^{K}(-1)\cdot\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}+\langle\frac{\partial g(\vec{I},p,\vec{D})}{\partial\vec{D}}|_{\vec{D}=\vec{D}_{t}(p)},\frac{\partial\vec{D}_{t}(p)}{\partial p}\rangle$
$=$ $-\sum_{k=1}^{K}\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}+\langle\frac{\partial g(\vec{I},p,\vec{D})}{\partial\vec{D}}|_{\vec{D}=\vec{D}_{t}(p)},-\vec{b}\rangle$
$\frac{\partial^{2}Q_{t}(\vec{I},p)}{\partial p^{2}}=$ $-2\sum_{k=1}^{K}\frac{\partial\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}}{\partial p}+\langle\mathcal{H}_{\vec{D}}g(\vec{I},p,\vec{D})^{\top}(-\vec{b}),-\vec{b}\rangle$
$=$ $-2\sum_{k=1}^{K}\frac{\partial\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}}{\partial p}+\vec{b}^{\top}\mathcal{H}_{\vec{D}}g(\vec{I},p,\vec{D})\vec{b}$
$=$ $2\langle\frac{\partial\sum_{k=1}^{K}-X_{i_{k},j_{k}}^{*}}{\partial\vec{D}},-\vec{b}\rangle+\vec{b}^{\top}\mathcal{H}_{\vec{D}}g(\vec{I},p,\vec{D})\vec{b},p\in[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}].$
Here $\mathcal{H}_{\vec{D}}g(\vec{I},p,\vec{D})$ represents the partial Hessian matrix of $g(\vec{I},p,\vec{D})$ with respect to $\vec{D}$. As we have shown in Lemma 3.2 that $g$ is marginally convex on $\vec{D}$, we know that $\mathcal{H}_{\vec{D}}g(\vec{I},p,\vec{D})\succeq 0$ and therefore $\vec{b}^{\top}\mathcal{H}_{\vec{D}}g(\vec{I},p,\vec{D})\vec{b}\geq 0$. Since $\vec{b}\geq 0$ as we have assumed, it is sufficient to show $\sum_{k=1}^{K}\sum_{k=1}^{K}-X_{i_{k},j_{k}}^{*}$ is monotonically non-increasing with respect to $\vec{D}$.
Now we Consider the following optimization problem
$R(\vec{I},\vec{D}):=\min_{X\in\mathbb{R}^{m\times n}}$ $\sum_{k=1}^{K}-\mathds{1}[\tilde{C}_{i_{k},j_{k}}<p]\cdot X_{i_{k},j_{k}}$ (22)
s.t. $X_{i,j}\geq 0$
$\sum_{i=1}^{m}X_{i,j}\leq D_{j},\forall j\in[n]$
$\sum_{j=1}^{n}X_{i,j}\leq I_{i},\forall i\in[m].$
According to Lemma A.5, we know that $R(\vec{I},\vec{D})$ is a convex function w.r.t. $[\vec{I};\vec{D}]$. Also, if we denote $\mathds{1}_{i_{k},j_{k}}:=\mathds{1}[\tilde{C}_{i_{k},j_{k}}<p]$, we will see $-\mathds{1}_{i_{k},j_{k}}\leq-\mathds{1}_{i_{k+1},j_{k+1}},\forall k\in[mn-1]$. This indicates the solution to Eq. 2 generated by Algorithm 1 is also an optimal solution to Eq. 22.
Consider an arbitrary $\Delta\vec{D}\geq\vec{0}$. Since the optimal solution $X^{*}$ to $R(\vec{I},\vec{D})$ is a feasible solution to $R(\vec{I},\vec{D}+\Delta\vec{D})$ (as the constraints get loose), we know that $R(\vec{I},\vec{D})\geq R(\vec{I},\vec{D}+\Delta\vec{D})$. Therefore, we have $\sum_{k=1}^{K}-X_{i_{k},j_{k}^{*}}$ is monotonically non-increasing with respect to $\vec{D}$. This holds the proof.
A.4 Proof of Lemma 3.5
We firstly propose and prove a lemma that shows the number of singularities of function $g$. Denote $\vec{I}^{*}(p):=\mathop{\mathrm{argmin}}_{\vec{I}}Q(\vec{I},p)$, and we have
Lemma A.6.
For any noise vector $\vec{N}_{t}\in\mathbb{R}^{n}$, there exists at most $3^{m+n}$ number of $p\in[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}]$ such that $g(\vec{I}^{*}(p),p,\vec{a}-\vec{b}\cdot p+\vec{N}_{t})$ is non-smooth with respect to $p$.
Proof of Lemma A.6.
Recall that $g(\vec{I},p,\vec{D}_{t}(p))=\sum_{k=1}^{K}(-p+C_{i_{k},j_{k}})\cdot X_{i_{k},j_{k}}^{*}$ where $X_{i_{k},j_{k}}=\min_{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}(p)}$. Therefore, the non-smoothness comes from the assignment of $X_{i_{k},j_{k}}$, which further comes from $\tilde{D}_{j_{k}}(p)$.
To facilitate the proof, we slightly enhance the lemma by letting $\vec{I}=\vec{u}+\vec{v}\cdot p$, and $\vec{D}=\vec{a}-\vec{b}\cdot p$ and prove the lemma for any $\vec{u},\vec{v},\vec{a},\vec{b}\geq 0$. (Here we merge $\vec{N}_{t}$ into $\vec{a}$ as it is fixed.)
We prove this lemma by induction on $m+n$.
  1. (i)
    When $m+n=2$ i.e. $m=n=1$, we have $g(\vec{I},p,\vec{D})=(-p+C_{1,1})\min\{u_{1}+v_{1}p,a_{1}-b_{1}p\}$, which has $1\leq 3^{m+n}$ singularities at most.
  2. (ii)
    Assume that the lemma holds for $\forall m,n,m+n\leq Z$ for some integer $Z\geq 2$. Now, consider the case when $m+n=Z+1$, and we have
    $\quad g(\vec{I},p,\vec{D})$ (23)
    $=(-p+C_{i_{1},j_{1}})\cdot\min\{u_{i_{1}}+v_{i_{1}}p,a_{j_{1}}-b_{j_{1}}\cdot p\}+\sum_{k=2}^{K}(-p+C_{i_{k},j_{k}})\cdot\min\{\tilde{I}_{i_{k}}(p),\tilde{D}_{j_{k}}(p)\}$
    $=\mathds{1}[u_{i_{1}}+v_{i_{1}}p\leq a_{j_{1}}-b_{j_{1}}\cdot p]\left((-p+C_{i_{1},j_{1}})(u_{i_{1}}+v_{i_{1}}p)+g(\vec{I}(p)-(u_{i_{1}}+v_{i_{1}}p)\vec{e}_{i_{1}},p,\vec{D}(p)-(u_{i_{1}}+v_{i_{1}}p)\vec{e}_{j_{1}})\right)$
    $\quad+\mathds{1}[u_{i_{1}}+v_{i_{1}}p>a_{j_{1}}-b_{j_{1}}\cdot p]\left((-p+C_{i_{1},j_{1}})(a_{j_{1}}-b_{j_{1}}p)+g(\vec{I}(p)-(a_{j_{1}}-b_{j_{1}}p)\vec{e}_{i_{1}},p,\vec{D}(p)-(a_{j_{1}}+b_{j_{1}}p)\vec{e}_{j_{1}})\right).$
    Here $\vec{e}_{i}:=[0,0,\ldots,0,1,0,\ldots,0]^{\top}$ is an indicator vector which has only one “$1$”. Notice that $g(\vec{I}(p)-(u_{i_{1}}+v_{i_{1}}p)\vec{e}_{i_{1}},p,\vec{D}(p)-(u_{i_{1}}+v_{i_{1}}p)\vec{e}_{j_{1}})$ and $g(\vec{I}(p)-(a_{j_{1}}-b_{j_{1}}p)\vec{e}_{i_{1}},p,\vec{D}(p)-(a_{j_{1}}+b_{j_{1}}p)\vec{e}_{j_{1}})$ is discussed in $m+n=Z$. Therefore, from the inductive hypothesis, we know that the total number of singularities is no more than $1+3^{Z}+3^{Z}\leq 3^{Z+1}$.
According to the principle of induction, we have proved this lemma valid. ∎

### v2

@@ 7 Conclusion
Appendix A Algorithmic Components: a Detailed Instructions
In what follows, we elaborate each component of $\mathcal{A}_{K}$’s algorithmic design in detail.
Horizontal search space for $p_{K}^{*}$. In the design of Stage $1$ algorithm as presented in Algorithm 3, we adopt the framework of zeroth-order online convex optimization. Specifically, we establish an epoch-based update rule of the search space of local optimal $p_{K}^{*}$. The search space (interval) for Epoch $\tau=1,2,\ldots$ is denoted as $[L_{K,\tau},U_{K,\tau}]$. Within each epoch, we divide the time horizon into a series of doubling sub-epochs to gather samples for $W(a),W(b),W(c)$ where $a,b,c$ are the three quarter points. By the end of each sub-epoch, we update the estimates and examine whether their estimation error bar is separable according to certain rules. As we keep doubling the size of sub-epochs, the estimation error bars are shrinking exponentially until they are separated. Then we reduce the search space by one quarter and proceed to Epoch $\tau+1$ Sub-Epoch $1$. When the search space is as sufficiently small as $O(1/T)$, we stop searching and proceed to Stage $2$.
Vertical uncertainty bound for $W(p_{K}^{*})$. In Stage $1$, we maintain an error bar $\Delta_{K}$ as the confidence bound of estimating each local optimal $W(p_{K}^{*})$. We show that the error bar has a size of $\tilde{O}(\frac{1}{\sqrt{T_{K}}})$ if we have run $\mathcal{A}_{K}$ for $T_{K}$ times so far. In addition to the statistical concentrations, another intuition of this fact comes from Lemma B.5: A not-distinguishable situation implies a comparable uncertainty bound for the optimal.
Complementary sampling to enhance $\Delta_{K}$. It is worth noting that $\Delta_{K}<+\infty$ does not exist for granted even after Stage 1. This is because we cannot update $\Delta_{K}$ when the search space $[L_{K},U_{K}]$ is updated, i.e., no simultaneous “horizontal converging” and “vertical converging”. As a consequence, if we are very “lucky” that we can always reduce the search space in the first sub-epoch of every epoch until $U_{K,\tau}-L_{K,\tau}\leq 1/T$, then we will have $\Delta_{K}=+\infty$ until Stage $2$. We resolve this issue in two approaches: (1) We upper bound the time periods before any $\Delta_{K}<+\infty$ by $O(\log T)$ based on the Pigeon-Hole Theorem. (2) We have Stage 2 as a complementary sampling stage without causing excessive regret. By the end of Stage $2$, we will have an ideal error bar for each agent $\mathcal{A}_{K}$.
Pure local exploitation contributing to global LCB. From Agent $\mathcal{A}_{K}$’s perspective, it runs pure exploitation in Stage 3 (Algorithm 5) without causing extra sub-regret. However, it still keeps updating the estimates of $\hat{W}_{K}$ and $\Delta_{K}$ to facilitate the LCB meta-algorithm.
Appendix B Proof Details
B.1 Proof of Lemma 3.1
@@ Appendix B Proof Details
$g(\vec{I},\vec{D}):=$ $\min_{x}\max_{\mu,\lambda,\eta\succeq 0}L(x,\mu,\lambda,\eta;\vec{I},\vec{D})$ (12)
$=$ $\max_{\mu,\lambda,\eta\succeq 0}\min_{x}L(x,\mu,\lambda,\eta;\vec{I},\vec{D})$
@@
B.2 Proof of Lemma 3.3
We denote $\tilde{C}_{i,j}(p):=-p+C_{i,j}$. For each fixed $k\in[mn]$, for any $p\in(C_{i_{k},j_{k}},C_{i_{k+1},j_{{k+1}}})$, the sign of any $\tilde{C}_{i,j}(p)$ is fixed. Therefore, when $k$ is fixed, we may let $X_{i,j}=0$ for any $(i,j)\succeq(i_{k+1},j_{k+1})$ in the optimal solution of $g(\vec{I},p,\vec{D})$. Given this, there exists $A\in\mathbb{R}^{k\times k},B\in\mathbb{R}^{k\times k},\vec{C}\in\mathbb{R}^{k}$ such that the optimization problem defined as Eq. 2 for $p\in(C_{i_{k},j_{k}},C_{i_{k+1},j_{{k+1}}})$ can be generalized to the following linear programming:
$g(\vec{I},p,\vec{D}):=$ $\min_{x\in\mathbb{R}^{k}}(-p\vec{\mathds{1}}+\vec{C})^{\top}x$ (13)
s.t. $x\succeq 0$
$Ax\preceq\vec{I}$
$Bx\preceq\vec{D}.$
Without loss of generality, in the following part of this proof of Lemma 3.3, we show that $g(\vec{I},p,\vec{D})$ defined in Eq. 13 is convex for $p\in(C_{i_{k},j_{k}},C_{i_{k+1},j_{{k+1}}})$. The Lagrangian of this new $g(\vec{I},p,\vec{D})$ is
$L(x,\mu,\lambda,\eta;\vec{I},\vec{D})$ (14)
::=:= $(-p\cdot\vec{\mathds{1}}+\vec{C})^{\top}x-\mu^{\top}x+\lambda^{\top}(Ax-\vec{I})+\eta^{\top}(Bx-\vec{D})$
$=$ $(-p\cdot\vec{\mathds{1}}+\vec{C}-\mu+A^{\top}\lambda+B^{\top}\eta)^{\top}x-\lambda^{\top}\vec{I}-\eta^{\top}\vec{D}.$
Since linear programming has strong duality, we further have
$g(\vec{I},p,\vec{D})=$ $\max_{\mu,\lambda,\eta}\min_{x}L(x,\mu,\lambda,\eta,\vec{I},\vec{D})$ (15)
$=$ $\max_{\mu,\lambda,\eta\geq 0}-\lambda^{\top}\vec{I}-\eta^{\top}\vec{D}$
[…]
As a consequence, the definition of $Q(\vec{I},p)$ is now generalized to
$Q_{t}(\vec{I},p):=$ $g(\vec{I},p,\vec{a}-\vec{b}p+\vec{N}_{t})$ (16)
$=$ $\max_{\lambda,\eta\geq 0}-\lambda^{\top}\vec{I}-\eta^{\top}(\vec{a}-\vec{b}p+\vec{N}_{t})$
[…]
Denote
$S(p):=\{(\lambda,\eta)|p\cdot\vec{\mathds{1}}\leq A^{\top}\lambda+B^{\top}\eta+\vec{C},\lambda\geq 0,\eta\geq 0\}.$ (17)
Notice that for any $p_{1},p_{2}$ s.t. $C_{i_{k},j_{k}}\leq p_{1}<p_{2}\leq C_{i_{k+1},j_{k+1}}$, we have $S(p_{1})\supseteq S(p_{2})$, indicating that $S(p)$ is a monotonically shrinking convex set as $p$ increases. Given the definition of $S(p)$ in Eq. 17, the definition of $Q_{t}(\vec{I},p)$ is equivalent to
$Q_{t}(\vec{I},p):=\max_{(\lambda,\eta)\in S(p)}-\lambda^{\top}\vec{I}-\eta^{\top}(\vec{a}-\vec{b}p+\vec{N}_{t})$ (18)
In the following, we prove a more generalized lemma, from which we can immediately derive the convexity of Eq. 18.
Lemma B.2.
Consider a family of functions :=$\mathcal{F}:=\{f(x;\theta)\}$, where $f(x;\theta)$ is Lipschitz and convex on $x\in[0,1]$ and is parametrized by $\theta$. $S(x)$ is a convex set that is monotonically not expanding w.r.t. $x$ (i.e., $S(x_{1})\supseteq S(x_{2})$ if $0\leq x_{1}<x_{2}\leq 1$). If $Q(x):=\max_{\theta\in S(x)}f(x;\theta)$ is Lipschitz, then $Q(x)$ is convex on $[0,1]$.
Proof of Lemma B.2.
Denote the epigraph of a function $g:\ \mathbb{R}\rightarrow\mathbb{R}$ as
$epi_{[l,r]}\ g(\cdot):=\{(x,t)|x\in[l,r],t\geq g(x)\}.$ (19)
Also, for set $S(x),x\in[0,1]$, denote
$S^{-1}(\theta):=\{x\in[0,1]|\theta\in S(x)\}$ (20)
Since $S(x)$ is monotonically not expanding, we know that $S^{-1}(\theta)$ is continuous. This is because $\theta\in S(x)$ is a sufficient condition of $\theta\in S(y)$ for any $y\in[0,x]$. Given these definitions, we have
$epi_{[0,1]}\ Q(\cdot)=$$\{(x,t)|x\in[0,1],t\geq Q(x)\}$ (21)
$=$ $\bigcup_{x\in[0,1]}\{(x,t)|t\geq\max_{\theta\in S(x)}f(x;\theta)\}$
$=$ $\bigcup_{x\in[0,1]}\{(x,t)|\forall\theta\in S(x),t\geq f(x;\theta)\}$
$=$ $\bigcup_{x\in[0,1]}\bigcap_{\theta\in S(x)}\{(x,t)|t\geq f(x;\theta)\}$
$=$ $\bigcap_{\theta\in S(0)}\bigcup_{x\in S^{-1}(\theta)}\{(x,t)|t\geq f(x;\theta)\}$
$=$ $\bigcap_{\theta\in S(0)}epi_{[0,\sup S^{-1}(\theta)]}\ f(\cdot;\theta).$
Since $f(x;\theta)$ is a convex function with respect to $x$ in $[0,1]$, we know that $epi_{[a,b]}f(\cdot;\theta)$ is a convex domain for any $0\leq a\leq b\leq 1$. Therefore, their intersections over all $\theta\in S(0)$ is a convex domain. This shows that the epigraph of $Q(x)$ is convex, and therefore $Q(x)$ is convex. ∎
With Lemma B.2, we know that $Q_{t}(\vec{I},p)$ defined in Eq. 18 is convex with respect to $p$. This ends the proof of Lemma 3.3.
B.3 Proof of Lemma 3.4